How many liters of chlorine at 101.325 kPa and 273K will be needed to make 75.0 grams of C2H2Cl4


1
Expert's answer
2021-05-17T04:17:12-0400

Using the ideal gas equation,


"PV= nRT"


"101325 \times V = \dfrac{75}{167.84}\times 8.31 \times 273"


"V = \dfrac{17006388}{170147.25}"


"V = 99.95L"